OpenStudy (anonymous):

You have 4 math books, 3 english books, and six science books. If you select books at random what is the probability of picking 3 english books?

OpenStudy (ganpat):

how many books u have ??

OpenStudy (ganpat):

@madison712 : lets do it together.. hope u dont mind..

OpenStudy (anonymous):

may be 13 :/

OpenStudy (ganpat):

add all them together.. you will get total number of books..

OpenStudy (ganpat):

you have total of 6+4+3 = 13 books.. now, picking book and it shd be English.. so u have 3 English books.. So the probability would be 13C3...

OpenStudy (anonymous):

even i think you are right :)

OpenStudy (amistre64):

probabilities aint greater than 1

OpenStudy (amistre64):

i think a hypergeometric probability might be more condusive

OpenStudy (amistre64):

\[\large\frac{\binom{M}{k}\binom{N-M}{n-k}}{\binom{N}{n}}\]

OpenStudy (ganpat):

oh sorry.. small mistake..

OpenStudy (ganpat):

it should be 3/13...

OpenStudy (ganpat):

@amistre64 : ?? what say ?

OpenStudy (amistre64):

or one way of looking at it might be if we put all 13 books into a box; the probability of picking an E in the first go around is: 3/13 out of the remaining 12 books ; the probability of picking an E in the 2nd go around is: 2/12 out of the remaining 11 books ; the probability of picking an E in the 3rd go around is: 1/11 P(A and B and C) = P(A) * P(B) * P(C)

OpenStudy (ganpat):

hmmm, agree.. i thought, just one book.. thnx !!!

OpenStudy (amistre64):

youre welcome